OpenStudy (johnnewman):

A jar contains 14 nickels, 10 dimes, 6 quarters, and 22 pennies. A coin is chosen at random from the jar. What is the probability that the coin chosen is a dime?

Mehek (mehek14):

\(\dfrac{dimes}{total}=\dfrac{10}{52}\) Simplify

OpenStudy (johnnewman):

i know 5 gose to ten two times

Mehek (mehek14):

Does 2 go into 52?

OpenStudy (johnnewman):

no

Mehek (mehek14):

Are you sure?

OpenStudy (johnnewman):

26

OpenStudy (johnnewman):

C.

Mehek (mehek14):

Correct

OpenStudy (johnnewman):

5/26

OpenStudy (johnnewman):

he data represent the number of phone calls received over an 8-day period. Semester Test, Part 1 What is the mean number of phone calls received? A. 6.6 B. 11 C. 5 D. 5.75

Mehek (mehek14):

\(\dfrac{4+5+9+1+0+5+11+11}{8}\)

OpenStudy (johnnewman):

D.

Mehek (mehek14):

Yes